OpenStudy (anonymous):

How do you graph a circle with center at (-1, -2), that also passes through the point (0, 2). What is the equation of the circle?

OpenStudy (anonymous):

Please help :c I have one day left to bump up my grade in geometry and I have no idea how to do this.

OpenStudy (campbell_st):

well the standard form is \[(x - h)^2 + (y - k)^2 = r^2\] (h, k) is the centre and r is the radius you were given the centre (-1, -2) so substitute them into the equation. \[(x + 1)^2 + (y + 2)^2 = r^2\] now to find the radius... one method is to use the distance formula and find the distance between the centre and the point (0, 2) another method is to substitute the point into the equation \[(x + 10^2 + (y + 2)^2 = r^2\] which will give you r^2 then you can get the equation

OpenStudy (anonymous):

Thank You!